# Reduced-Order Hydrodynamic Modelling of a Sphere Near a Wall Using Sparse Regression and Neural Operators

### v1

@@ Introduction
Conventional tools for predicting ship motions, such as strip theory and linear potential-flow methods, are well established and computationally efficient under the assumptions of inviscid, irrotational flow and small-amplitude motions [Price1974]. In this linear regime, hydrodynamic forces are commonly decomposed into linear added mass, radiation damping, hydrostatic stiffness, and first-order wave-excitation components [WAMIT, Han2024]. However, as wave steepness, motion amplitude, and geometric complexity increase, this linear description becomes progressively less accurate: higher-order wave interactions, non-linear wave–radiation, and non-linear hydrostatic effects are no longer adequately captured [Sulisz1993, IEAOES2024].Similar limitations have been reported for ship–ship interactions in close proximity, where wave resonance in gaps and strong reflection effects introduce additional non-linearities [Tan2017, Zhou2023b].
To resolve these effects with higher fidelity, recent studies have turned to Computational Fluid Dynamics (CFD), which can represent viscous effects, and non-linear free-surface dynamics at the expense of substantially increased computational cost [Kramer2021].
However, because high-fidelity CFD remains too costly for real-time use, recent work in naval hydrodynamics has increasingly turned to surrogate modelling strategies capable of reproducing key nonlinear behaviours at a fraction of the computational expense. Data-driven sequence models including LSTM [Hao2022, Liong2024, Tang2021, Xu2021, Sun2022], and hybrid CNN–RNN architectures [Cheng2021, Li2022, Zhang2019, Shi2025] have been increasingly applied to predict ship motions in waves showing some promising results for accurate real-time short-term motion prediction [DAgostino2022, Silva2022]. Modal-decomposition approaches such as POD and DMD have also been used to construct low-dimensional representations that can be combined with neural networks for nonlinear forecasting [Francesco2021, Diez2024]. More recently, physics-informed (PINNs) [Schirmann2023, Guan2024] and operator-learning methods (Onets) [Zhao2025], have been introduced to wave–body interaction problems to improve generalisation in data-sparse regimes. In parallel, sparse-regression techniques such as SINDy have shown that compact and interpretable nonlinear ODEs can be recovered directly from time series dynamics data [Brunton2016, Huang2023] .
Despite this progress, most existing surrogate models focus on isolated single-body dynamics. Comparatively little work has addressed the multi-scale, multi-body interactions characteristic of launch-and-recovery operations [Casey2019, Zhou2023b], motivating the need for physics-interpretable surrogates such as the one developed in this study.
@@
In previous work, the present authors developed and validated a CFD framework for a simplified LARS scenario in which a small USV is idealised as a one-degree-of-freedom (1-DoF) heaving sphere operating near a large vertical wall representing the mothership shown in Figure 1 and 2. This idealised problem is chosen because it retains the key hydrodynamic mechanisms relevant to launch and recovery while remaining well-supported by prior studies, including experimental measurements of sphere heave-decay motion [Kramer2021], responses of a sphere under regular wave forcing [IEAOES2024], and analytically derived solutions [Hulme1981].
@@
Previous work by this team [Hoffman2025] have demonstrated that CFD can capture key nonlinear hydrodynamic through comparing the simulation output with benchmark experimental data and analytical solutions. Based on the validated flow model, a parametric dataset of heave-decay responses across wall distances $WD\in(9,25)$m and drop heights $DH\in(0.5,5)$m was constructed. The drop heights considered are intentionally kept less than the sphere’s radius to ignore physics relating to an entry-type problem. The heave response dataset, depicted in Figure 3, is normalised by $\frac{Hv}{WD}$. A strong sensitivity on the input conditions (WD,DH) is evident.
@@
Surface tension effects are neglected, justified by a Weber number well in excess of unity and under-relaxation factors for momentum, viscosity, and body forces are reduced $(n<1)$ to stabilise convergence. Two additional equations are solved for capturing the turbulence in the flow by employing the $k-\omega$ SST turbulence model. A turbulence damping source term was added to limit the over prediction of viscosity near the water-air interface, which arises from the VOF formulation whereby the turbulence models tend to over predict eddy viscosity from the high strain rate at the interface [ANSYS2025].
@@
The linear added mass $A$ and radiation damping $B$ coefficients for the heaving sphere are obtained analytically in the work of [Hulme1981].
@@
Figure 4: Validation of drop test from drop height of (a) $0.1$ and (b) $0.5$ of the spheres diameter. Comparison between our CFD simulation, experimental data [Kramer2021] and linear model.
@@
where $A_{i}$ is the incident wave amplitude, and $\omega$ is the frequency of the incident waves. $X$ is the frequency dependent excitation force coefficient of a sphere, calculated using the commercial potential flow solver WAMIT in the work by [Kramer2021].
@@
Given analytically derived linear coefficients, the following linear model for the RAO is given in the paper by [IEAOES2024]:
@@
Figure 6 shows a comparison between the linear model and the CFD derived RAOs. It is evident that the CFD results systematically depart from the linear model (Equation 14) when the wave period is close to the spheres fundamental frequency and as wave steepness increase. This behaviour is consistent with the observations of [Kramer2021], who showed that higher-order hydrodynamic effects become increasingly significant with larger wave steepness $(s)$. Correspondingly, the PSD in Figure 6 reveals a redistribution of energy from the fundamental frequency $\omega$ into higher harmonics (e.g. $2\omega$ and $3\omega$). As noted by [Schoukens2019], a linear time-invariant system cannot transfer energy between frequencies; thus, the appearance of these harmonics provides clear evidence of nonlinear dynamical behaviour.
@@
For launch-and-recovery scenarios several effects undermine linear assumptions. Large relative motions, steep waves, and geometric complexity introduce; higher-order wave interactions, non-linear changes in hydrostatic restoring force due to the instantaneous submerged volume, and non-linear excitation forces. Previous CFD and experimental studies of heaving bodies in waves have shown that these mechanisms lead to significant deviations from linear theory [IEAOES2024, Sulisz1993], wherby excitation and radiation forces are no longer separable by simple superposition.
@@
Sparsity is then enforced via Sequential Thresholded Least Squares (STLSQ) [Huang2023]. At each iteration $k$:
@@
SINDy identifies sinusoidal excitation terms whose phase aligns closely with the wave-travel-time prediction. Around WD=15 m, a minimal excitation forces amplitude is observed (Figure 15), which is the same distance where the wave is 90 degrees offset from the spheres motion. No simple analytical relation for the excitation force magnitude has been found, but would likely require a similar methodology as [Li2021].

### v2

@@ Introduction
Conventional tools for predicting ship motions, such as strip theory and linear potential-flow methods, are well established and computationally efficient under the assumptions of inviscid, irrotational flow and small-amplitude motions [19]. In this linear regime, hydrodynamic forces are commonly decomposed into linear added mass, radiation damping, hydrostatic stiffness, and first-order wave-excitation components [28, 9]. However, as wave steepness, motion amplitude, and geometric complexity increase, this linear description becomes progressively less accurate: higher-order wave interactions, non-linear wave–radiation, and non-linear hydrostatic effects are no longer adequately captured [24, 18].Similar limitations have been reported for ship–ship interactions in close proximity, where wave resonance in gaps and strong reflection effects introduce additional non-linearities [26, 32].
To resolve these effects with higher fidelity, recent studies have turned to Computational Fluid Dynamics (CFD), which can represent viscous effects, and non-linear free-surface dynamics at the expense of substantially increased computational cost [14].
However, because high-fidelity CFD remains too costly for real-time use, recent work in naval hydrodynamics has increasingly turned to surrogate modelling strategies capable of reproducing key nonlinear behaviours at a fraction of the computational expense. Data-driven sequence models including LSTM [10, 17, 27, 29, 25], and hybrid CNN–RNN architectures [5, 16, 30, 22] have been increasingly applied to predict ship motions in waves showing some promising results for accurate real-time short-term motion prediction [6, 23]. Modal-decomposition approaches such as POD and DMD have also been used to construct low-dimensional representations that can be combined with neural networks for nonlinear forecasting [1, 7]. More recently, physics-informed (PINNs) [20, 8] and operator-learning methods (Onets) [31], have been introduced to wave–body interaction problems to improve generalisation in data-sparse regimes. In parallel, sparse-regression techniques such as SINDy have shown that compact and interpretable nonlinear ODEs can be recovered directly from time series dynamics data [3, 12] .
Despite this progress, most existing surrogate models focus on isolated single-body dynamics. Comparatively little work has addressed the multi-scale, multi-body interactions characteristic of launch-and-recovery operations [4, 32], motivating the need for physics-interpretable surrogates such as the one developed in this study.
@@
In previous work, the present authors developed and validated a CFD framework for a simplified LARS scenario in which a small USV is idealised as a one-degree-of-freedom (1-DoF) heaving sphere operating near a large vertical wall representing the mothership shown in Figure 1 and 2. This idealised problem is chosen because it retains the key hydrodynamic mechanisms relevant to launch and recovery while remaining well-supported by prior studies, including experimental measurements of sphere heave-decay motion [14], responses of a sphere under regular wave forcing [18], and analytically derived solutions [13].
@@
Previous work by this team [11] have demonstrated that CFD can capture key nonlinear hydrodynamic through comparing the simulation output with benchmark experimental data and analytical solutions. Based on the validated flow model, a parametric dataset of heave-decay responses across wall distances $WD\in(9,25)$m and drop heights $DH\in(0.5,5)$m was constructed. The drop heights considered are intentionally kept less than the sphere’s radius to ignore physics relating to an entry-type problem. The heave response dataset, depicted in Figure 3, is normalised by $\frac{Hv}{WD}$. A strong sensitivity on the input conditions (WD,DH) is evident.
@@
Surface tension effects are neglected, justified by a Weber number well in excess of unity and under-relaxation factors for momentum, viscosity, and body forces are reduced $(n<1)$ to stabilise convergence. Two additional equations are solved for capturing the turbulence in the flow by employing the $k-\omega$ SST turbulence model. A turbulence damping source term was added to limit the over prediction of viscosity near the water-air interface, which arises from the VOF formulation whereby the turbulence models tend to over predict eddy viscosity from the high strain rate at the interface [2].
@@
The linear added mass $A$ and radiation damping $B$ coefficients for the heaving sphere are obtained analytically in the work of [13].
@@
Figure 4: Validation of drop test from drop height of (a) $0.1$ and (b) $0.5$ of the spheres diameter. Comparison between our CFD simulation, experimental data [14] and linear model.
@@
where $A_{i}$ is the incident wave amplitude, and $\omega$ is the frequency of the incident waves. $X$ is the frequency dependent excitation force coefficient of a sphere, calculated using the commercial potential flow solver WAMIT in the work by [14].
@@
Given analytically derived linear coefficients, the following linear model for the RAO is given in the paper by [18]:
@@
Figure 6 shows a comparison between the linear model and the CFD derived RAOs. It is evident that the CFD results systematically depart from the linear model (Equation 14) when the wave period is close to the spheres fundamental frequency and as wave steepness increase. This behaviour is consistent with the observations of [14], who showed that higher-order hydrodynamic effects become increasingly significant with larger wave steepness $(s)$. Correspondingly, the PSD in Figure 6 reveals a redistribution of energy from the fundamental frequency $\omega$ into higher harmonics (e.g. $2\omega$ and $3\omega$). As noted by [21], a linear time-invariant system cannot transfer energy between frequencies; thus, the appearance of these harmonics provides clear evidence of nonlinear dynamical behaviour.
@@
For launch-and-recovery scenarios several effects undermine linear assumptions. Large relative motions, steep waves, and geometric complexity introduce; higher-order wave interactions, non-linear changes in hydrostatic restoring force due to the instantaneous submerged volume, and non-linear excitation forces. Previous CFD and experimental studies of heaving bodies in waves have shown that these mechanisms lead to significant deviations from linear theory [18, 24], wherby excitation and radiation forces are no longer separable by simple superposition.
@@
Sparsity is then enforced via Sequential Thresholded Least Squares (STLSQ) [12]. At each iteration $k$:
@@
SINDy identifies sinusoidal excitation terms whose phase aligns closely with the wave-travel-time prediction. Around WD=15 m, a minimal excitation forces amplitude is observed (Figure 15), which is the same distance where the wave is 90 degrees offset from the spheres motion. No simple analytical relation for the excitation force magnitude has been found, but would likely require a similar methodology as [15].